OpenStudy (anonymous):

MEDAL GIVEN TO CORRECT ANSWER. If a racehorse starts from rest and accelerates at a rate of 4.7 m/s2, how long does it take the horse to go 20 m? A. 4.3 s B. 2.9 s C. 8.5 s D. 9.1 s

OpenStudy (anonymous):

s = distance = (1/2) (acceleration) (time)^2 time = sqrt [2 (distance)/(acceleration)] t = 2.9 s